OpenStudy (anonymous):

What is the equation of the line that passes through (5, -2) and (-3, 4)? 3x - 4y - 7 = 0 3x + 4y - 7 = 0 -3x + 4y - 7 = 0

OpenStudy (anonymous):

help plzzz

OpenStudy (anonymous):

um let me think ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@welshfella @Compassionate

OpenStudy (snowflake251):

y = 7/4-(3 x)/4

OpenStudy (snowflake251):

x-intercept | 7/3~~2.33333 y-intercept | 7/4 = 1.75 slope | -3/4 = -0.75

OpenStudy (photon336):

@taivane easiest way to do this is to find the slope of those two points.

OpenStudy (photon336):

The idea is that if you find the slope, it would be clear that the line will pass through those two points. \[y = m(x)+b \] \[m = \frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\] \[\frac{ -2-(4)}{ 5-(-3) } = \frac{ -6 }{ 8 } = \frac{ -3 }{ 4 } = slope,m \]

OpenStudy (compassionate):

\[y - y _{1} = m(x - x_{1})\] Points are (5, -2) and (-3, 4)